40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 19 18:56:55 2018, Jon, Update, General, AUX beam alignment issues 
    Reply  Wed Jun 20 00:57:04 2018, johannes, Update, General, AUX beam alignment issues as110_rf_30MHz.pdfyend_rf_30MHz.pdf
       Reply  Sun Jun 24 23:13:46 2018, johannes, Update, General, AUX beam alignment issues AS_AUX_SETUP.pdf
Message ID: 13989     Entry time: Wed Jun 20 00:57:04 2018     In reply to: 13987     Reply to this: 14013
Author: johannes 
Type: Update 
Category: General 
Subject: AUX beam alignment issues 

We did swap a lens as discussed in elog 13968, but they both had f=100mm specified, the difference being one was AR-coated for 1064 and bi-convex, while the other one was plano-convex and had a different coating. The reason for the large beam spot was something else: The fiber wasn't sitting in the coupler properly. When reconnecting the fiber after taking it out make sure to align the key on the fiber end with the notch in the coupler before tightening. After discovering this the following was done:

  • Fixed fiber mounting situation
  • Tested AUX alignment into fiber on PSL table, was still good
  • The AUX polarization was aligned to the wrong fiber axis. I fixed this. The coupler on the PSL table has it's noth oriented vertically since we're using s-polarized light. The AS-table coupler is rotated by 90 degrees, such that the notch points to the side. This way we technically don't need any halfwaveplates for rotation. However, there are still current HWPs installed.
  • Locked both arms and ran dither alignment until satisfactory
  • Misaligned ITMX and ETMX, and further set the ITMX pitch offset to 0.0
  • Started overlapping the expectedly misaligned beams by eye. For this I turned the power of the deflected beam down to 50mV bias voltage, which gives the PSL and AUX lasers similar card-brightness on the shared path
  • Misaligned SRM more because there was still the strong prompt reflection coming out the AS port.
  • Restored phaselock between AUX and PSL, with beat at 30MHz between 1st-order diffracted in fiber and PSL
  • Immediately saw STRONG 30MHz RF signal on AG4395. Disappeared when blocking AUX, and optimized alignment brought the signal up to -10dBm, as shown in attachment #1
  • Checked YEND PDA10CF and saw a -80dBm RF signal at 30 MHz (#2), compatible with earlier observations.

Before leaving I restored the XARM alignment. SRM remains misaligned, LSC off. Alignment shouldn't change drastically over night, so I suggest when picking this work up tomorrow to directly look for the beats after phaselocking AUX and PSL

Attachment 1: as110_rf_30MHz.pdf  23 kB  Uploaded Wed Jun 20 02:56:51 2018  | Hide | Hide all
as110_rf_30MHz.pdf
Attachment 2: yend_rf_30MHz.pdf  23 kB  Uploaded Wed Jun 20 03:08:46 2018  | Hide | Hide all
yend_rf_30MHz.pdf
ELOG V3.1.3-